# Generation of mechanical cat-like states via optomagnomechanics

### v1

@@ Abstract
We propose an optomagnomechanical approach for preparing a cat-like superposition state of mechanical motion. Our protocol consists of two steps and is based on the magnomechanical system where the magnetostrictively induced displacement further couples to an optical cavity mode via radiation pressure. We first prepare a squeezed mechanical state by driving the magnomechanical system with a two-tone microwave field. We then switch off the microwave drives and send a weak red-detuned optical pulse to the optical cavity to weakly activate the optomechanical anti-Stokes scattering. We show that $k$ phonons can be subtracted from the prepared squeezed state, conditioned on the detection of $k$ anti-Stokes photons from the cavity output field, which prepares the mechanical motion in a cat-like state. The work provides a new avenue for preparing mechanical superposition states by combining opto- and magnomechanics and may find applications in the study of macroscopic quantum states and the test of collapse theories.
@@ I Introduction
Here, we present a scheme for preparing a mechanical cat-like state based on the OMM system. We take the approach of subtracting excitations from the squeezed vacuum state MD97 ; AB07 . Specifically, our scheme consists of two steps: $i$) Preparing a mechanical squeezed state in the magnomechanical subsystem by applying a two-tone driving field to the magnon mode; $ii$) Subtracting phonons from the generated squeezed state by sending a weak red-detuned optical pulse to the optomechanical cavity. The detection of $k$ anti-Stokes photons in the cavity output field heralds a $k$-phonon-subtracted squeezed state, corresponding to a cat-like state of mechanical motion.
The paper is organized as follows. In Sec. II, we introduce the OMM system and show how to generate a mechanical squeezed state in the magnomechanical subsystem. In Sec. III, we show how to subtract $k$ phonons from the mechanical squeezed state in the optomechanical subsystem via the detection of the scattered anti-Stokes photons, and present the results of the mechanical cat-like states. Finally, we summarize our findings in Sec. IV.
II Magnomechanical SQUEEZING OF MECHANICAL motion
Figure 1: (a) Two-step protocol adopted to generate cat-like states of mechanical motion in an OMM system. First step: the magnon mode is driven by two microwave fields at the frequencies $\omega_{\pm}=\omega_{m}\pm\omega_{b}$ to prepare a mechanical squeezed state. Second step: a weak red-detuned optical pulse is sent to the optomechanical cavity to subtract phonons, conditioned on the detection of anti-Stokes photons in the cavity output field. (b) Frequencies of the magnomechanical subsystem and two microwave driving fields in the first step, and of the optomechanical subsystem and the weak optical pulse in the second step.
@@ II Magnomechanical SQUEEZING OF MECHANICAL motion
In the first step, we aim to squeeze the mechanical motion by driving the magnon with two microwave fields (Fig. 1(b)). Due to the absence of a strong laser driving field to enhance the optomechanical coupling, the effective optomechanical coupling is very weak. The optical cavity is essentially decoupled from the driven magnomechanical system. Therefore, the tripartite OMM system reduces to an effective two-mode system with the Hamiltonian given by
@@
which indicates that the magnon mode is driven by two microwave fields at the frequencies $\omega_{\pm}=\omega_{m}\pm\omega_{b}$, and $\Omega_{\pm}$ denote the associated magnon-drive coupling strength, i.e., the Rabi frequency JL18 .
@@
Here, $G_{\pm}=G_{0}m_{\pm}$ are the enhanced magnomechanical coupling strengths due to the two strong driving fields at frequencies $\omega_{\pm}$, which can be set real by adjusting the phases of the drive fields to have real $m_{\pm}$. Due to the time-dependent fast oscillating terms in Eq. (5), the above QLEs are difficult to solve. However, under the condition of $\kappa_{b},\kappa_{m},G_{\pm}\ll\omega_{b}$ JL15 , the rotating-wave approximation can be made by neglecting the fast oscillating terms. Consequently, we obtain
@@
Figure 2: Wigner function of the squeezed mechanical mode. The dashed circle corresponds to vacuum fluctuation. The parameters are provided in the main text.
Equation (6) can be rewritten in terms of the quadrature fluctuations, which can be cast in the following matrix form
@@
Owing to the linearized dynamics and Gaussian input noises, the steady state of the system is a two-mode Gaussian state, which is fully characterized by a $4\times 4$ covariance matrix (CM) $V$, with its entries defined as $V_{ij}=\langle u_{i}(t)u_{j}(t)+u_{j}(t)u_{i}(t)\rangle/2$ ($i,j=1,2$). The steady-state CM can be obtained by directly solving the Lyapunov equation DV07
$\begin{split}AV+VA^{T}=-D,\end{split}$ (9)
@@
Figure 3: (a) Degree of squeezing $S$ (dB) of the mechanical mode versus $G_{+}/G_{-}$ for various temperatures. We fix the coupling $G_{-}/2\pi=0.1$ MHz and vary $G_{+}$. (b) Degree of squeezing $S$ (dB) versus bath temperature $T$ and the coupling $G_{-}$. The coupling $G_{+}$ is optimized for each value of $G_{-}$, and the other parameters are as those in Fig. 2.
In Fig. 3(a), we plot the degree of squeezing $S$ (in units of dB) as a function of the ratio $G_{+}/G_{-}$ for various bath temperatures, and $S$ is defined as $S=-10{\rm log}10[V_{\rm min}/V_{\rm vac}]$, with $V_{\rm min}$ being the minimum variance of the mechanical quadrature and $V_{\rm vac}$ being the variance of the vacuum fluctuation. Clearly, there is an optimal ratio $G_{+}/G_{-}$, which decreases as the temperature rises. This is due to the fact that, on the one hand, the difference between the couplings $G_{+}$ and $G_{-}$ should be sufficiently large in order to efficiently cool the mechanical mode; on the other hand, almost equal couplings $G_{+}\simeq G_{-}$ is required to have an extremely large squeezing Jie17 . The trade-off between these two opposite trends results in an optimal ratio $G_{+}/G_{-}$. As the temperature rises, a larger cooling rate is needed to cool the mechanical mode with more thermal noise, which thereby leads to a reduced optimal ratio and a reduced degree of squeezing. In Fig. 3(b), we plot the degree of squeezing $S$ versus the coupling strength $G_{-}$ and the temperature $T$. It tells that the maximum degree of squeezing increases with the coupling $G_{-}$ (with $G_{+}$ being optimized) and decreases with the temperature, in consistency with the results of Fig. 3(a) for a fixed $G_{-}$.
III PHONON SUBTRACTION via optical pulses
Once the mechanical mode is prepared in a squeezed state (ideally a squeezed vacuum state, but practically a squeezed thermal state due to the added thermal noise), we switch off the two microwave drive fields and then send an optical pulse to the optomechanical subsystem to implement the phonon subtraction operation, which we show below can yield a mechanical cat-like state.
Since the magnon driving fields are turned off, the effective magnomechanical coupling becomes very weak. The system, under an optical driving field, then becomes an effective two-mode system with an optomechanical Hamiltonian, given by
@@ III PHONON SUBTRACTION via optical pulses
Similarly, we linearize the dynamics following the procedures in Sec. II. When the drive field is red-detuned from the cavity by $\Delta\equiv\omega_{c}-\omega_{0}=\omega_{b}$, and in the resolved sideband limit $\omega_{b}\gg\kappa_{c}$, we derive the following linearized QLEs for the quantum fluctuations $\{\delta c$, $\delta b\}$, which, in the interaction picture with respect to $\hbar\omega_{0}c^{\dagger}c$, are given by JL21
@@
where $c_{in}$ is the input noise of the cavity, and the effective optomechanical coupling $G_{c}=g_{0}\langle c\rangle$, with $\langle c\rangle=\frac{iE}{i\Delta+\frac{\kappa_{c}}{2}}\simeq\frac{E}{\omega_{b}}$. For simplicity, we consider a flattop pulse such that $G_{c}$ is constant during the pulse. Note that in getting Eq. (13), we have neglected the frequency shift induced by the optomechanical interaction as it is typically much smaller than $\Delta=\omega_{b}$.
We consider the pulse duration to be much shorter than the mechanical lifetime, such that the mechanical dissipation is negligible during the pulse. The corresponding QLEs, in the interaction picture with respect to $\hbar\omega_{c}c^{\dagger}c+\hbar\omega_{b}b^{\dagger}b$, are given by
@@
We use weak pulses, yielding $G_{c}\ll\kappa_{c}$, which allows us to adiabatically eliminate the cavity, and obtain $\delta c=i\frac{2G_{c}}{\kappa_{c}}\delta b+\frac{2}{\sqrt{\kappa_{c}}}c_{in}$. By using the input-output relation $c_{out}=\sqrt{\kappa_{c}}\delta c-c_{in}$, we get
@@
$\begin{split}C_{out}(t)&=e^{-Gt}C_{in}(t)+i\sqrt{1-e^{-2Gt}}b(0),\\ \delta b(t)&=e^{-Gt}\delta b(0)+i\sqrt{1-e^{-2Gt}}C_{in}(t).\end{split}$ (17)
@@
with $C_{n}=\left(\frac{e^{i\phi}}{2}\right)^{n}\frac{\sqrt{(2n)!}}{n!}$. The function of the pulse is equivalent to applying the propagator $U(t)$ onto the initial state $|0\rangle_{c}|\xi\rangle_{b}$, where $|0\rangle_{c}$ denotes the cavity is in vacuum. At the end of the pulse, the system evolves to be
$\begin{split}|\psi\rangle_{c,b}&=U|0\rangle_{c}|\xi\rangle_{b}=\sum_{n=0}^{\infty}\frac{(i\tan\theta)^{\textit{n}}}{n!}\left(C_{in}^{\dagger}b\right)^{n}|0\rangle_{c}|\xi^{\prime}\rangle_{b}\\ &\approx|0\rangle_{c}|\xi^{\prime}\rangle_{b}+i\tan\theta|1\rangle_{c}\big(\textit{b}|\xi^{\prime}\rangle_{\textit{b}}\big)-\frac{\tan^{2}\theta}{\sqrt{2}}|2\rangle_{\textit{c}}\left(\textit{b}^{2}|\xi^{\prime}\rangle_{\textit{b}}\right),\end{split}$ (20)
@@
In getting the last line of Eq. (20), we have omitted higher-order terms with $n>2$, which is a good approximation when $\tan^{2}\theta\ll 1$, i.e., $Gt\ll 1$. Equation (20) indicates that the subtraction of a single phonon $b|\xi^{\prime}\rangle_{b}$ can be realized when a single anti-Stokes photon in the cavity output field is detected, and the successful probability is approximately $\tan^{2}\theta$. Similarly, the subtraction of two phonons $b^{2}|\xi^{\prime}\rangle_{b}$ can also be achieved, with a much smaller probability $\sim\tan^{4}\theta$, conditioned on the detection of a two-photon state. We note that due to the pulse interaction, the squeezing $r$ of the squeezed vacuum state is slightly reduced, because $\tanh r\rightarrow\tanh r\,\cos^{2}\theta$ as seen in Eq. (21).
In a practical situation, as demonstrated in Sec. II, we prepare a squeezed thermal state due to the nonnegligible thermal noise at a finite temperature, which is reflected in the CM $V_{b}$ of the mechanical mode. To be compatible with the propagator $U(t)$, we transform the CM into a density matrix $\rho_{in,b}$ (Appendix), which acts as the initial state before the optical pulse is applied. Combining the vacuum state of the cavity, the initial state of the optomechanical system is then $\rho_{in}=\rho_{in,b}\otimes|0\rangle\langle 0|_{c}$, which, at the end of the pulse, evolves into the state $\rho_{b,c}=U\rho_{in}U^{\dagger}$, given by MD97 ; MB94
@@
$\begin{split}\rho_{b,c}&=\sum^{\infty}_{n=0}\frac{(\tan\theta)^{2n}}{n!}b^{n}|\cos\theta|^{b^{\dagger}b}\rho_{in,b}|\cos\theta|^{b^{\dagger}b}b^{\dagger n}\otimes|\textit{n}\rangle\langle\textit{n}|_{\textit{c}}\\ &\approx|\cos\theta|^{b^{\dagger}b}\rho_{in,b}|\cos\theta|^{b^{\dagger}b}\otimes|0\rangle\langle 0|_{\textit{c}}\\ &+\tan^{2}\theta\,b|\cos\theta|^{b^{\dagger}b}\rho_{in,b}|\cos\theta|^{b^{\dagger}b}b^{\dagger}\otimes|1\rangle\langle 1|_{\textit{c}}\\ &+\frac{\tan^{4}\theta}{2}b^{2}|\cos\theta|^{b^{\dagger}b}\rho_{in,b}|\cos\theta|^{b^{\dagger}b}b^{\dagger 2}\otimes|2\rangle\langle 2|_{\textit{c}},\end{split}$ (23)
@@
where higher-order terms with $n>2$ are neglected because $\tan^{2}\theta\ll 1$. The above equation indicates that a single phonon (two phonons) can be subtracted from the initial squeezed thermal state, i.e., $b\rho_{in,b}b^{\dagger}$ ($b^{2}\rho_{in,b}b^{\dagger 2}$), conditioned on the detection of a single-photon (two-photon) state in the cavity output. Note that the slightly reduced $r$ as revealed in Eq. (21) is not visible because a general mixed state $\rho_{in,b}$ is adopted in Eq. (23).
@@
Figure 4: Wigner function of (a) the single-phonon [(b) the two-phonon] subtracted squeezed mechanical state. The parameters are provided in the text.

### v2

@@ Abstract
We propose an optomagnomechanical approach for preparing a cat-like superposition state of mechanical motion. Our protocol consists of two steps and is based on the magnomechanical system where the magnetostrictively induced displacement further couples to an optical cavity mode via radiation pressure. We first prepare a squeezed mechanical state by driving the magnomechanical system with two microwave pulses. We then switch off the microwave drives and send a weak red-detuned optical pulse to the optical cavity to weakly activate the optomechanical anti-Stokes scattering. We show that $k$ phonons can be subtracted from the prepared squeezed thermal state, conditioned on the detection of $k$ anti-Stokes photons from the cavity output field, which prepares the mechanical motion in a cat-like state. The work provides a new avenue for preparing mechanical superposition states by combining opto- and magnomechanics and may find applications in the study of macroscopic quantum states and the test of collapse theories.
@@ I Introduction
Here, we present a scheme for preparing a mechanical cat-like state based on the OMM system using fast microwave and optical pulses. We adopt the approach of subtracting excitations from the squeezed vacuum state MD97 ; AB07 ; TJ16 . Specifically, our scheme consists of two steps: $i$) Preparing a mechanical squeezed state in the magnomechanical subsystem by applying a pair of red- and blue-detuned microwave pulses to the magnon mode; $ii$) Subtracting phonons from the generated squeezed state by sending a weak red-detuned optical pulse to the optomechanical cavity. The detection of $k$ anti-Stokes photons in the cavity output field heralds a $k$-phonon-subtracted squeezed state, corresponding to a cat-like state of mechanical motion.
The paper is organized as follows. In Sec. II, we introduce the OMM system and the pulse sequence used in our protocol. We then show how to generate a mechanical squeezed state in the magnomechanical subsystem by applying two microwave pulses in Sec. III, and how to subtract $k$ phonons from the mechanical squeezed state in the optomechanical subsystem by applying a red-detuned optical pulse in Sec. IV. Finally, we summarize our findings in Sec. V.
II The OMM system and the pulse sequence
Figure 1: (a) Two-step protocol adopted to generate cat-like states of mechanical motion in an OMM system. First step: the magnon mode is driven by two microwave pulses at frequencies $\omega_{\pm}=\omega_{m}\pm\omega_{b}$ to prepare a mechanical squeezed state. Second step: a weak red-detuned optical pulse is sent to the optomechanical cavity to subtract phonons, conditioned on the detection of anti-Stokes photons in the cavity output field. (b) The pulse sequence used in our protocol. Two microwave pulses at frequencies $\omega_{m}\pm\omega_{b}$ and with duration $\tau_{\rm sq}$ are used to prepare a mechanical squeezed state. A red-detuned optical pulse with duration $\tau_{s}$ is sent to the cavity to subtract phonons from the squeezed state, conditioned on the detection of anti-Stokes photons. Once a cat-like state is generated, another red-detuned optical pulse with duration $\tau_{r}$ is sent to read out the mechanical state for subsequent tomography operation. (c) Frequencies of the magnomechanical subsystem and two microwave driving fields in the first step, and of the optomechanical subsystem and the weak optical pulse in the second step.
@@ II The OMM system and the pulse sequence
Our protocol adopts a series of fast microwave and optical pulses, as shown in Fig. 1(b). In the first step, two microwave pulses with different frequencies and duration $\tau_{\rm sq}$ are used to squeeze the mechanical motion Schwab15 . Once a squeezed state is prepared and after a short period when magnons die out, a red-detuned optical pulse with duration $\tau_{s}$ is sent to the optical cavity to subtract phonons from the squeezed state. After anti-Stokes photons being detected from the cavity output field, which heralds a mechanical cat-like state, another red-detuned optical pulse with duration $\tau_{r}$ is sent to map the mechanical state to the cavity output field Jie18a ; Tan20 for subsequent tomography of the state.
III Magnomechanical SQUEEZING OF MECHANICAL motion
In the first step, to simplify the model we consider two flattop microwave pulses Jie18a . During the pulse, the driving strength can be assumed constant, which corresponds to a constant Rabi frequency introduced below. After a short period $\tau_{\rm sq}$, when the mechanical system evolves into a stationary state with an invariant degree of squeezing, we then switch off the microwave pulses. Due to the absence of a strong laser driving field to enhance the optomechanical coupling, the effective optomechanical coupling is very weak. The optical cavity is essentially decoupled from the driven magnomechanical system. Therefore, the tripartite OMM system reduces to an effective two-mode system with the Hamiltonian given by
@@ III Magnomechanical SQUEEZING OF MECHANICAL motion
which indicates that the magnon mode is driven by two microwave fields at the frequencies $\omega_{\pm}=\omega_{m}\pm\omega_{b}$, and $\Omega_{\pm}$ denote the associated magnon-drive coupling strength, i.e., the Rabi frequency. For a YIG micro bridge, the Rabi frequency is defined as $\Omega=\frac{\sqrt{5}}{4}\gamma\sqrt{N}B_{0}$ JL18 , where the gyromagnetic ratio $\gamma/2\pi=28$ GHz/T, the total number of spins $N=\rho V$ with $V$ being the volume of the micro bridge and $\rho=4.22\times 10^{27}$m-3 being the spin density of the YIG, and $B_{0}$ is the amplitude of the driving magnetic field. The drive amplitude is related to the drive power via $B_{0}=\sqrt{2\mu_{0}P_{0}/(lwc)}$ QST23 , where $\mu_{0}$ is the vacuum magnetic permeability, $c$ is the speed of the electromagnetic wave propagating in vacuum, and $l$ and $w$ are the length and width of the micro bridge, which is approximated as a cube for simplicity.
@@
Here, $G_{\pm}=G_{0}m_{\pm}$ are the enhanced magnomechanical coupling strengths due to the two strong driving fields at frequencies $\omega_{\pm}$, which can be set real by adjusting the phases of the drive fields to have real $m_{\pm}$. Due to the time-dependent fast oscillating terms in Eq. (5), the above QLEs are difficult to solve. However, under the condition of $\kappa_{b},\kappa_{m},G_{\pm}\ll\omega_{b}$ JL15 , the rotating-wave approximation (RWA) can be made by neglecting the fast oscillating terms. Consequently, we obtain
@@
Eq. (6) can be rewritten in terms of the quadrature fluctuations, which can be cast in the following matrix form
@@
Owing to the linearized dynamics and Gaussian input noises, the state of the system at any time is a two-mode Gaussian state, which is fully characterized by a $4\times 4$ covariance matrix (CM) $V$, with its entries defined as $V_{ij}=\langle u_{i}(t)u_{j}(t)+u_{j}(t)u_{i}(t)\rangle/2$ ($i,j=1,2$). The dynamical CM $V$ can be obtained by solving the equation
$\begin{split}\dot{V}=AV+VA^{T}+D,\end{split}$ (9)
@@
Figure 2: (a) Degree of squeezing (dB) of the mechanical mode versus time. The blue curve is obtained by using Eq. (6) with the RWA, while the thinner red curve is achieved using the full equations Eq. (5) without the RWA. (b) Wigner function of the squeezed mechanical mode. The dashed circle corresponds to vacuum fluctuation. The parameters are provided in the main text.
@@
Figure 3: (a) Degree of squeezing (dB) of the mechanical mode versus $G_{+}/G_{-}$ for various temperatures. We fix the coupling $G_{-}/2\pi=0.1$ MHz and vary $G_{+}$. (b) Degree of squeezing $S$ (dB) versus bath temperature $T$ and the coupling $G_{-}$. The coupling $G_{+}$ is optimized for each value of $G_{-}$, and the other parameters are as those in Fig. 2.
@@
Figure 4: Degree of squeezing (dB) of the mechanical mode versus mechanical damping rate $\kappa_{b}$. The other parameters are as those in Fig. 2
In Fig. 3(a), we plot the degree of squeezing $S$ as a function of the ratio $G_{+}/G_{-}$ for various bath temperatures. Clearly, there is an optimal ratio $G_{+}/G_{-}$, which decreases as the temperature rises. This is due to the fact that, on the one hand, the difference between the couplings $G_{+}$ and $G_{-}$ should be sufficiently large in order to efficiently cool the mechanical mode; on the other hand, almost equal couplings $G_{+}\simeq G_{-}$ is required to have an extremely large squeezing Jie17 . The trade-off between these two opposite trends results in an optimal ratio $G_{+}/G_{-}$. As the temperature rises, a larger cooling rate is needed to cool the mechanical mode with more thermal noise, which thereby leads to a reduced optimal ratio and a reduced degree of squeezing. In Fig. 3(b), we plot the degree of squeezing $S$ versus the coupling strength $G_{-}$ and the temperature $T$. It tells that the maximum degree of squeezing increases with the coupling $G_{-}$ (with $G_{+}$ being optimized) and decreases with the temperature, in consistency with the results of Fig. 3(a) for a fixed $G_{-}$. Note that the microwave drive pulses might cause heating of the YIG crystal. Nevertheless, the possible heating leads to a temperature rise, whose effect is included in Fig. 3(b). It shows that even for the temperature being up to hundreds of mK, the degree of squeezing is still considerable. In Fig. 4, we further analyze the effect of the mechanical damping rate. Clearly, the degree of squeezing remains high, about 7 dB, for $\kappa_{b}$ being up to 1 kHz, corresponding to a mechanical $Q$ factor of $3\times 10^{4}$.
IV PHONON SUBTRACTION via optical pulses
Once the mechanical mode is prepared in a squeezed state (ideally a squeezed vacuum state, but practically a squeezed thermal state due to the added thermal noise IS20 ; Tan20 ; Schwab15 ), we switch off the two microwave pulses. After a short period for magnons to decay, we then send an optical pulse to the optomechanical subsystem to implement the phonon subtraction operation, which we show below can yield a mechanical cat-like state.
Since the microwave driving fields are turned off, the effective magnomechanical coupling becomes very weak. The system, under an optical driving field, then becomes an effective two-mode system with an optomechanical Hamiltonian, given by
@@ IV PHONON SUBTRACTION via optical pulses
Similarly, we linearize the dynamics following the procedures in Sec. III. When the drive field is red-detuned from the cavity by $\Delta\equiv\omega_{c}-\omega_{0}=\omega_{b}$, and in the resolved sideband limit $\omega_{b}\gg\kappa_{c}$, we derive the following linearized QLEs for the quantum fluctuations $\{\delta c$, $\delta b\}$, which, in the interaction picture with respect to $\hbar\omega_{0}c^{\dagger}c$, are given by JL21
@@
where $c_{in}$ is the input noise of the cavity, and the effective optomechanical coupling $G_{c}=g_{0}\langle c\rangle$, with $\langle c\rangle=\frac{iE}{i\Delta+\frac{\kappa_{c}}{2}}\simeq\frac{E}{\omega_{b}}$. Again, we consider a flattop pulse such that $G_{c}$ is constant during the pulse. Note that in getting Eq. (13), we have neglected the frequency shift induced by the optomechanical interaction as it is typically much smaller than $\Delta=\omega_{b}$.
We consider the pulse duration to be much shorter than the mechanical life and coherence time, i.e., $\tau_{s}\ll(N_{b}\kappa_{b})^{-1}\simeq 0.25$ ms, such that the mechanical dissipation and decoherence are negligible during the pulse IS20 ; Tan20 ; Jie18a . The corresponding QLEs, in the interaction picture with respect to $\hbar\omega_{c}c^{\dagger}c+\hbar\omega_{b}b^{\dagger}b$, are given by
@@
The pulse is sufficiently weak, yielding $G_{c}\ll\kappa_{c}$, which allows us to adiabatically eliminate the cavity, and obtain $\delta c=i\frac{2G_{c}}{\kappa_{c}}\delta b+\frac{2}{\sqrt{\kappa_{c}}}c_{in}$. By using the input-output relation $c_{out}=\sqrt{\kappa_{c}}\delta c-c_{in}$, we get
@@
$\begin{split}C_{out}(t)&=e^{-Gt}C_{in}(t)+i\sqrt{1-e^{-2Gt}}\delta b(0),\\ \delta b(t)&=e^{-Gt}\delta b(0)+i\sqrt{1-e^{-2Gt}}C_{in}(t).\end{split}$ (17)
@@
with $C_{n}=\left(\frac{e^{i\phi}}{2}\right)^{n}\frac{\sqrt{(2n)!}}{n!}$. As will be shown later, the analysis of this ideal squeezed vacuum pure state is important, because it allows us to see more clearly how the squeezing parameter $r$ evolves due to the pulse interaction. The function of the pulse is equivalent to applying the propagator $U(t)$ onto the initial state $|0\rangle_{c}|\xi\rangle_{b}$, where $|0\rangle_{c}$ denotes the cavity is in vacuum. At the end of the pulse, the system evolves to be
$\begin{split}|\psi\rangle_{c,b}&=U|0\rangle_{c}|\xi\rangle_{b}=\sum_{n=0}^{\infty}\frac{(i\tan\theta)^{\textit{n}}}{n!}\left(C_{in}^{\dagger}b\right)^{n}|0\rangle_{c}|\xi^{\prime}\rangle_{b}\\ &\approx|0\rangle_{c}|\xi^{\prime}\rangle_{b}+i\tan\theta|1\rangle_{c}\left(\textit{b}|\xi^{\prime}\rangle_{\textit{b}}\right)-\frac{\tan^{2}\theta}{\sqrt{2}}|2\rangle_{\textit{c}}\left(\textit{b}^{2}|\xi^{\prime}\rangle_{\textit{b}}\right)\\ &{-i\frac{\tan^{3}\theta}{\sqrt{6}}|3\rangle_{c}\left(b^{3}|\xi^{\prime}\rangle_{b}\right)},\end{split}$ (20)
@@
In getting the last line of Eq. (20), we have omitted higher-order terms with $n>2$, which is a good approximation when $\tan^{2}\theta\ll 1$, i.e., $Gt\ll 1$. Equation (20) indicates that the subtraction of a single phonon $b|\xi^{\prime}\rangle_{b}$ can be realized when a single anti-Stokes photon in the cavity output field is detected, and the successful probability is approximately $\tan^{2}\theta$. Similarly, the subtraction of two (three) phonons $b^{2}|\xi^{\prime}\rangle_{b}$ ($b^{3}|\xi^{\prime}\rangle_{b}$) can also be achieved, with a much smaller probability $\sim\tan^{4}\theta$ ($\tan^{6}\theta$), conditioned on the detection of a two (three)-photon state. We note that due to the pulse interaction, the squeezing $r$ of the squeezed vacuum state is slightly reduced, because $\tanh r\rightarrow\tanh r\,\cos^{2}\theta$ as seen in Eq. (21).
In a practical situation, we prepare a squeezed thermal mixed state, which is reflected in the CM $V_{b}$ of the mechanical mode obtained in Sec. III, from which two key characteristic parameters of the squeezed thermal state are extracted, i.e., the squeezing parameter $r=1.23$, and the residual phonon number $\bar{n}=0.014$ (see Appendix for details). To be compatible with the propagator $U(t)$, we transform the CM $V_{b}$ into a density matrix $\rho_{in,b}$ (Appendix), which acts as the initial state before the optical pulse is applied. Combining the vacuum state of the cavity, the initial state of the optomechanical system is then $\rho_{in}=\rho_{in,b}\otimes|0\rangle\langle 0|_{c}$, which, at the end of the pulse, evolves into the state $\rho_{b,c}=U\rho_{in}U^{\dagger}$, given by MD97 ; MB94
@@
$\begin{split}\rho_{b,c}&=\sum^{\infty}_{n=0}\frac{(\tan\theta)^{2n}}{n!}b^{n}|\cos\theta|^{b^{\dagger}b}\rho_{in,b}|\cos\theta|^{b^{\dagger}b}b^{\dagger n}\otimes|\textit{n}\rangle\langle\textit{n}|_{\textit{c}}\\ &\approx|\cos\theta|^{b^{\dagger}b}\rho_{in,b}|\cos\theta|^{b^{\dagger}b}\otimes|0\rangle\langle 0|_{\textit{c}}\\ &+\tan^{2}\theta\,b|\cos\theta|^{b^{\dagger}b}\rho_{in,b}|\cos\theta|^{b^{\dagger}b}b^{\dagger}\otimes|1\rangle\langle 1|_{\textit{c}}\\ &+\frac{\tan^{4}\theta}{2}b^{2}|\cos\theta|^{b^{\dagger}b}\rho_{in,b}|\cos\theta|^{b^{\dagger}b}b^{\dagger 2}\otimes|2\rangle\langle 2|_{\textit{c}}\\ &+{\frac{\tan^{6}\theta}{6}b^{3}|\cos\theta|^{b^{\dagger}b}\rho_{in,b}|\cos\theta|^{b^{\dagger}b}b^{\dagger 3}\otimes|3\rangle\langle 3|_{\textit{c}}},\end{split}$ (23)
@@
where higher-order terms with $n>3$ are neglected because of their extremely small probabilities. The above equation indicates that $k$-phonon can be subtracted from the initial squeezed thermal state, i.e., $b^{k}\rho_{in,b}b^{\dagger k}$ ($k=1,2,3$), conditioned on the detection of a $k$-photon state in the cavity output. Note that the slightly reduced $r$ revealed in Eq. (21) is not visible because a general mixed state $\rho_{in,b}$ is adopted in Eq. (23).
@@
Figure 5: Wigner function of (a) the single-phonon (b) the two-phonon (c) the three-phonon subtracted squeezed mechanical state. The parameters are provided in the text.